QUESTION 33

  1. Rollins Corporation's bonds have a 12 percent coupon, paid semiannually. It matures in 20 years, and its current market price is $1,000. The maturity value is $1,000. The firm's marginal tax rate is 40 percent. Which of the following is Rollins' after-tax cost of debt?

    A.

    6.8 percent

    B.

    7.2 percent

    C.

    5.3 percent

    D.

    9.5 percent

    E.

    8.4 percent
Answer #1

As the bond is trading at par, YTM and coupon rate will be same.

Using the formula for calculating YTM,

YTM =(Coupon + (Face value - Price)/n)/ (Face value +Price )/2

Coupon = 12% of 1000 = 120. The semi-annual coupon is 60

Face value = Price = 1000.

n = 20*2 = 40 periods.

YTM = (60+ (1000-1000)/40))/(1000+1000)/2 = 60/1000 = 0.06 or 6%

The annual YTM is 6 % * 2 = 12%.

After tax cost of debt = 12*(1-tax rate) =12*(1-0.40)= 12*0.60 = 7.2%